Find f1(x) if f(x) = 3x - 4 이 0
Answer #1

f(x) = 3x - 4

Let y = f(x) implies x = f-1(y)

Now

y = f(x)

y = 3x - 4

y + 4 = 3x

(y + 4)/3 = x

x = (y + 4)/3

implies

f-1(y) = (y + 4)/3

Therefore

f-1(x) = (x + 4)/3 i.e., second option is correct